# Go Beyond Earth: Understanding Human Actions and Scenes in Microgravity Environments

### v1

@@ 1 Introduction
Figure 1: An illustration of the MicroG-4M dataset, which contains videos from real and simulated microgravity environments (e.g., movies). The dataset supports benchmarks for three tasks: (1) video captioning, (2) video-based question answering, and (3) fine-grained human action recognition under microgravity.
@@
Yuri Gagarin’s historic flight in 1961 marked the beginning of human space exploration. Since then, significant milestones have been achieved, including crewed lunar landings, the continuous operation of the International Space Station (ISS) for over $25$ years, and the participation of more than $650$ individuals in space missions [1]. With numerous planned crewed missions in the near future, the frequency and complexity of human activities in space are expected to increase substantially. In this context, ensuring the safety, enhancing the operational efficiency of space missions, and safeguarding the health and well-being of astronauts are of paramount importance.
@@
With the rapid advancement of artificial intelligence, the integration of robotic systems aboard spacecraft is anticipated in the near future. These systems will assist astronauts in routine and mission-critical tasks. Consequently, there is a growing demand for the development of deep learning-based scene understanding and action recognition methods tailored specifically for the unique challenges posed by the microgravity environment. Human action recognition [2, 3, 4], scenario captioning [5], and Visual Question Answering (VQA) [6] are essential for intelligent human-robot collaboration, particularly in space, where precise perception and understanding of astronauts’ actions and their surrounding context are crucial for ensuring operational safety, efficiency, and autonomous assistance under constrained conditions. This capability is crucial for ensuring mission efficiency, enhancing astronaut safety, and providing autonomous assistance in the confined and complex conditions of space habitats. Numerous human action recognition datasets and video caption datasets have been developed, playing an important role in various research areas within computer vision and in industrial applications [7, 8, 9, 10].
@@
However, most of the existing datasets for video scenario captioning and action recognition are recorded on Earth without microgravity settings, e.g., Kinetics400 [11], AVA [12], FineGym [13], and ActivityNet Captions [14].
Human actions in space differ significantly from those on Earth due to microgravity, where basic movements like standing, walking, eating, and object manipulation follow entirely different physical principles [15]. For instance, ‘standing’ can occur at any angle, ‘walking’ resembles swimming, and objects are often grabbed mid-air rather than picked up or put down conventionally. As a result, existing Earth-based datasets fail to capture these nuances, limiting the effectiveness of Human Action Recognition (HAR) in microgravity environments.
@@
To address this gap, we introduce MicroG-4M, a new video benchmark specifically designed for spatio-temporal and semantic understanding of human activities in microgravity. The name “4M” reflects four core characteristics of the dataset: Multi-domain, Multimodal, Multi-task, and Microgravity. MicroG-4M comprises $4,759$ three-second video clips drawn from public YouTube footage of real space missions and carefully selected, realistic space-themed films to augment scenario diversity and coverage. The clips span scenes inside the ISS, the Tiangong Space Station, crewed spacecraft cabins, and extravehicular activities. The corpus contains more than $390,000$ annotated frames with bounding boxes for each visible individual and $13,000+$ action labels covering $50$ distinct actions. In addition, it includes $1,238$ descriptive clip captions created by human annotators and $7,000+$ open-ended question–answer pairs aimed at assessing factual, causal, and counterfactual understanding of scenes in micro-gravity. Both captions and QA annotations were carefully curated through multiple refinement rounds and data selection stages, ensuring high semantic fidelity, contextual accuracy, and relevance to the microgravity setting. Together, these resources enable multi-label spatio-temporal detection, fine-grained action recognition, caption generation, and VQA within a single dataset. To evaluate existing methods, we build the first comprehensive benchmark for these tasks, dubbed MicroG-Bench. We evaluated representative video encoders (SlowFast [4], X3D [3], and MViT [16]) for human action recognition and leading vision-language models (e.g., InternVideo [17], Gemini 1.5 Pro [18], GPT-4o [19]) for captioning and VQA tasks. Across all evaluated subtasks, state-of-the-art terrestrial models experience significant performance degradation, underscoring the unique challenges posed by micro-gravity environments, such as arbitrary orientations, floating objects, and confined spacecraft interiors. These results highlight the necessity of dedicated benchmarks to facilitate the advancement of more robust and generalizable AI systems tailored for space applications. MicroG-Bench, therefore, provides a unified, fair yardstick that will guide the development of more robust and generalizable perception and language systems for astronaut assistance and autonomous mission operations. Then we present a short summarization of the contribution in this work. We present MicroG-4M, the first video-based dataset for activity recognition and scene understanding specifically designed for the micro-gravity environment. Additionally, we establish the first benchmark to evaluate state-of-the-art human action recognition methods and vision-language models for fine-grained action recognition and video captioning/QA in this unique setting.
@@ 2 Related Work
Vision-Based Understanding in Microgravity and Space Environments. Microgravity challenges core assumptions of terrestrial vision systems, necessitating perception models tailored to space habitats. Early work focused on visual localization, where traditional SLAM proved unreliable, prompting robust alternatives using visual-inertial fusion, semantic mapping, and CAD-informed constraints [20, 21, 22, 23], validated on platforms like Astrobee [24]. Real-time vision has also enabled dynamic tasks such as target tracking and scene change detection [25, 26].
As research shifted toward interaction, studies explored astronaut pose recovery [27, 28], gesture recognition [29, 30], EMG-based input [31], and long-term astronaut identification [32]. However, high-level semantic understanding—such as reasoning about astronaut actions and intent—remains limited. Current assistants like CIMON [33] offer basic perception but lack contextual reasoning. To fill this gap, we introduce a unified benchmark for action localization, video captioning, and visual question answering in operational space environments.
@@
Datasets for Action Detection, Video Captioning, and VQA. Progress in video understanding has been largely enabled by the introduction of benchmark datasets across action detection, captioning, and VQA. For action detection, early datasets such as UCF101 [34] and HMDB51 [35] provided trimmed classification tasks, later extended by Kinetics [11] and ActivityNet [36] to large-scale, untrimmed settings with diverse action categories. AVA [12] further introduced spatio-temporal annotations for atomic actions, enabling fine-grained multi-label detection in realistic scenes. Video captioning evolved from MSVD [37] and MSR-VTT [38], which paired short clips with multiple sentences, to dense captioning in untrimmed videos via ActivityNet Captions [39] and procedural datasets like YouCook2 [40]. Multilingual benchmarks such as VATEX [41] expanded the task to cross-lingual settings with high-quality parallel annotations. In VQA, static image datasets like VQA v2.0 [42] and CLEVR [43] laid the foundation for compositional reasoning, while TGIF-QA [44], MovieQA [45], and TVQA [46] introduced temporal and multimodal reasoning in video. Recent efforts such as NExT-QA [47] and CLEVRER [48] focus on causal inference and counterfactual reasoning. However, existing datasets are constrained to terrestrial environments and lack domain-specific complexity inherent to space-based activities. To bridge this gap, we introduce the first benchmark for video captioning and VQA in microgravity, enabling the evaluation of vision-language models in long-horizon, space-relevant scenarios.
@@
Fine-Grained Human Action Recognition. Fine-grained video-based action recognition [49, 50] aims to detect fine-grained, indivisible human actions in both single- and multi-person videos. Unlike standard clip-level recognition, atomic action localization requires frame-level, multi-label classification with spatio-temporal bounding box predictions. To tackle this, CNN-based [2, 3, 4] and transformer-based [51, 2, 16, 52, 17, 53, 49] models have been adapted by adding multi-label heads, bounding box regressors, and region-of-interest modules. Notably, Ryali et al.[51] introduced a hierarchical vision transformer balancing accuracy and efficiency, while Wanget al. [2] proposed a dual-masked autoencoder for improved video pretraining. In this work, we evaluate such HAR methods for the first time in microgravity scenarios.
@@ 3 Collection Methodology
Filtering. Video clips undergo automatic filtering based on person detection and scene-transition analysis. Specifically, we employed YOLOv11 [54] for human detection and PySceneDetect [55] to identify abrupt scene changes, discarding clips with insufficient human-action content or disrupted temporal continuity.
Bounding-box Annotation. Person bounding boxes are automatically annotated using YOLOv11 [54] detection combined with BoT-SORT tracking [56]. To enhance annotation accuracy, adaptive strategies were employed by assessing video motion intensity using sparse optical flow methods [57, 58, 59]. Annotation parameters were dynamically adjusted accordingly to optimize computational efficiency and annotation precision. The resulting structured annotations include bounding-box coordinates, unique identities, and detection confidence scores.
@@
Action Label Annotation. In this phase, we established a specialized action taxonomy tailored to microgravity environments, adapted from the AVA dataset’s original 80 atomic actions [12]. To ensure environmental applicability and semantic consistency, we systematically excluded irrelevant terrestrial or water-specific actions, merged closely related actions into unified categories, and introduced contextually appropriate semantic adjustments. Furthermore, clear differentiation criteria were defined for visually or semantically similar actions, thereby standardizing the annotation protocol. Subsequently, each three-second video clip was manually annotated as a discrete unit, assigning up to five distinct visible or inferable action labels per detected individual. The finalized annotations were exported as structured CSV files for subsequent utilization.
@@
For Video Question Answering (VQA), we adopted a structured approach based on [60], formulating diverse questions using standard interrogative forms. Based on the complexity of the video captions, Large Language Models (LLMs) generate a proportional number of candidate question-answer pairs, explicitly filtering out cases with missing or uncertain information. These candidates are then ranked by the LLMs according to logical consistency, linguistic fluency, semantic relevance, and informational value. Subsequently, multiple rounds of human-assisted review and prompt adjustment are employed to eliminate any hallucinated or fabricated content produced by the LLMs. Finally, a concise set of six diverse and content-rich QA pairs is selected for each video, ensuring comprehensive coverage from broad contextual understanding to detailed actions.
@@ 4 Dataset Composition
Fine-Grained Human Action Recognition Subset: Contains fine-grained, multi-label annotations for spatiotemporal human action detection, comprising $4,759$ manually annotated three-second video clips from authentic microgravity environments. Each clip includes annotations for up to five distinct actions per detected individual. The $50$ action labels are organized into three categories: Object Manipulation ($4,976$ annotations, $37.60\%$), Person Interaction ($1,198$ annotations, $32.34\%$), and Person Movement ($1,037$ annotations, $30.07\%$). Figure 2(d) illustrates the distribution of these categories in both real and simulated video clips, while Figure 2(e) shows the distribution of person counts per clip, highlighting the diversity of social configurations. Overall, the subset contains $13,261$ action annotations, including $9,610$ from real footage, $3,651$ from simulated sources, and $390,000$ bounding box annotations.
@@
Figure 2 presents representative video clips from the dataset, demonstrating typical annotation examples and highlighting the visual and semantic diversity of the dataset. More details regarding the dataset can be found in supplementary.
@@ 5.1 Benchmark Protocol
Data Split. We partition the dataset into training, validation, and test subsets in a 7:1:2 ratio. Row-level distribution is as follows: Training set: 9,266 records (69.93% of 13,251); Validation set: 1,329 records (10.03% of 13,251); Test set: 2,656 records (20.04% of 13,251). Video-level distribution is as follows: Training set: 3,331 videos (69.99% of 4,759); Validation set: 475 videos (9.98% of 4,759); Test set: 953 videos (20.03% of 4,759).
Baselines for Fine-Grained HAR in Micro-Gravity Scenarios. For fine-grained action recognition, we evaluate well-established baselines from video-based human action recognition, including transformer-based models (MViTv1 [61], MViTv2 [16]) and CNN-based models (I3D [62], SlowFast [4], X3D [3], C2D [4]). These widely used architectures cover both paradigms and allow us to assess generalization to the unique spatiotemporal dynamics of microgravity. All models are initialized with Kinetics400 [11] pretrained weights for better convergence.
Baselines methods for Video Captioning and QA in Micro-Gravity Scenarios. For video captioning and question answering in microgravity, we evaluate strong baselines, including open-source models (VideoChatGPT [63], mPLUG-Owl-3 [64], LLaVA-Next [65], VideoLLaVa [66], Qwen-2.5-VL [67], InternVideo [17]) and closed-source models (ChatGPT-4o [19], Gemini 1.5 Pro [68]). Open-source models offer reproducibility via public weights and code, while closed-source models serve as upper-bound references. This mix enables both transparent analysis and comprehensive evaluation of video-language understanding in microgravity.
Evaluation Metrics for Fine-Grained HAR. Our evaluation metrics include mAP@0.5, F1 score, recall, and AUROC, all calculated using the macro method. Among these, mAP@0.5 is the primary metric for measuring average detection accuracy per category and thus comprehensively evaluating the model’s action recognition performance in a microgravity environment.
Evaluation Metrics for Video Caption and QA. We adopt standard automatic evaluation metrics, including CIDEr [69], BLEU-4 [70], ROUGE-L [71], METEOR [72], and BERTScore (F1) [73], all rescaled to a 0$\sim$100 range for consistency. For semantic similarity, we report S-BERT [74] and S-VQA [75] scores, both computed as cosine similarity between Sentence-BERT [76] embeddings of predicted and reference texts. S-VQA is used specifically for answer evaluation in generative visual question answering settings, capturing semantic equivalence beyond lexical overlap. More details regarding the implementations can be found in the supplementary materials.

### v3

@@ 1 Introduction
Yuri Gagarin’s historic flight in 1961 marked the beginning of human space exploration. Since then, significant milestones have been achieved, including crewed lunar landings, the continuous operation of the International Space Station (ISS) for over $25$ years, and the participation of more than $650$ individuals in space missions Moskowitz & Wolf (2025). With numerous planned crewed missions in the near future, the frequency and complexity of human activities in space are expected to increase substantially. In this context, ensuring the safety, enhancing the operational efficiency of space missions, and safeguarding the health and well-being of astronauts are of paramount importance.
@@
With the rapid advancement of artificial intelligence, the integration of robotic systems aboard spacecraft is anticipated in the near future. These systems will assist astronauts in routine and mission-critical tasks. Consequently, there is a growing demand for the development of deep learning-based scene understanding and action recognition methods tailored specifically for the unique challenges posed by the microgravity environment.
Human action recognition Wang et al. (2023a); Feichtenhofer (2020); Feichtenhofer et al. (2019), scenario captioning Chen et al. (2024), and Visual Question Answering (VQA) Hu et al. (2023) are essential for intelligent human-robot collaboration, particularly in space, where precise perception and understanding of astronauts’ actions and their surrounding context are crucial for ensuring operational safety, efficiency, and autonomous assistance under constrained conditions. This capability is crucial for ensuring mission efficiency, enhancing astronaut safety, and providing autonomous assistance in the confined and complex conditions of space habitats. Numerous human action recognition datasets and video caption datasets have been developed, playing an important role in various research areas within computer vision and in industrial applications Camarena et al. (2023); Pareek & Thakkar (2021); Al-Faris et al. (2020); Le et al. (2022).
@@
However, most of the existing datasets for video scenario captioning and action recognition are recorded on Earth without microgravity settings, e.g., Kinetics400 Kay et al. (2017), AVA Gu et al. (2018), FineGym Shao et al. (2020), and ActivityNet Captions Chen et al. (2019).
Human actions in space depart markedly from their terrestrial counterparts because microgravity removes gravity-aligned orientation and reliable support surfaces. Basic behaviors—standing, locomotion, eating, and manual manipulation—follow different kinematics and contact patterns Hagio et al. (2022). Standing becomes orientation-invariant, often maintained via foot restraints or handholds rather than ground support; locomotion is achieved by drifting or pulling along structures rather than gait; and manipulation frequently involves releasing or catching free-floating objects instead of placing or picking from a surface. These shifts violate terrestrial orientation, support/contact, and object-dynamics priors, which helps explain the degradation of Earth-trained Human Action Recognition (HAR) models in orbit and motivates a microgravity-specific benchmark.
Figure 1: An illustration of the MicroG-4M, containing videos from real and simulated microgravity environments (e.g., movies). The dataset supports benchmarks for three tasks: (1) video captioning, (2) video question answering, and (3) fine-grained human action recognition under microgravity.
To address this gap, we introduce MicroG-4M, a new video benchmark specifically designed for spatio-temporal and semantic understanding of human activities in microgravity. The name “4M” reflects four characteristics: Multi-source (Real mission footage and physically plausible film), Multimodal (RGB + text annotations), Multi-task (HAR, captioning, VQA), and Microgravity.
@@
MicroG-4M comprises $4,759$ three-second video clips drawn from public YouTube footage of real space missions and carefully selected, realistic space-themed films to augment scenario diversity and coverage. The clips span scenes inside the ISS, the Tiangong Space Station, crewed spacecraft cabins, and extravehicular activities. The corpus contains more than $390,000$ annotated frames with bounding boxes for each visible individual and $13,000+$ action labels covering $50$ distinct actions. In addition, it includes $1,238$ descriptive clip captions created by human annotators and $7,000+$ open-ended question–answer pairs aimed at assessing factual, causal, and counterfactual understanding of scenes in microgravity. Both captions and QA annotations were carefully curated through multiple refinement rounds and data selection stages, ensuring high semantic fidelity, contextual accuracy, and relevance to the microgravity setting. Together, these resources enable multi-label spatio-temporal detection, fine-grained action recognition, caption generation, and VQA within a single dataset. To evaluate existing methods, we build the first comprehensive benchmark for these tasks, dubbed MicroG-Bench. We evaluate representative video encoders (SlowFast Feichtenhofer et al. (2019), X3D Feichtenhofer (2020), and MViT Li et al. (2022)) for human action recognition and leading vision-language models (e.g., InternVideo Wang et al. (2022), Gemini 1.5 Pro Anil et al. (2023), GPT-4o Hurst et al. (2024)) for captioning and VQA tasks. Across all evaluated subtasks, state-of-the-art terrestrial models experience significant performance degradation, underscoring the unique challenges posed by microgravity environments, such as arbitrary orientations, floating objects, and confined spacecraft interiors. These results highlight the necessity of dedicated benchmarks to facilitate the advancement of more robust and generalizable AI systems tailored for space applications. MicroG-Bench, therefore, provides a unified, fair yardstick that will guide the development of more robust and generalizable perception and language systems for astronaut assistance and autonomous mission operations. We summarize our contributions as follows. We present MicroG-4M, the first video-based dataset for activity recognition and scene understanding specifically designed for the microgravity environment. Additionally, we establish the first benchmark to evaluate state-of-the-art human action recognition methods and vision-language models for fine-grained action recognition and video captioning/QA in this unique setting.
@@ 2 Related Work
Vision-Based Understanding in Microgravity and Space Environments. Microgravity challenges terrestrial vision assumptions, requiring perception models for space habitats. Early work showed traditional SLAM to be unreliable, leading to robust alternatives with visual-inertial fusion, semantic mapping, CAD-informed constraints Soussan et al. (2022); Mao et al. (2024); Miller et al. (2022); Tweddle et al. (2015), and validated on platforms like Astrobee Kang et al. (2024). Real-time vision has also enabled dynamic tasks such as target tracking and scene change detection Oestreich et al. (2021); Dinkel et al. (2024). Research has shifted toward interaction, including astronaut pose recovery Gan et al. (2023); Ouyang et al. (2025), gesture recognition Lingyun et al. (2020); Gao et al. (2020), and EMG-based input Assad et al. (2013). Yet, high-level semantic understanding of astronaut actions and intent remains limited. Current assistants like CIMON Eisenberg et al. (2024) provide basic perception but lack contextual reasoning. To address this, we propose a unified benchmark for action recognition, video captioning, and visual question answering in space operations.
@@
Datasets for Action Detection, Video Captioning, and VQA. Progress in video understanding has been largely enabled by the introduction of benchmark datasets across action detection, captioning, and VQA. For action detection, early datasets such as UCF101 Soomro et al. (2012) and HMDB51 Kuehne et al. (2011) provided trimmed classification tasks, later extended by Kinetics Kay et al. (2017) and ActivityNet Caba Heilbron et al. (2015) to large-scale, untrimmed settings with diverse action categories. AVA Gu et al. (2018) further introduced spatio-temporal annotations for atomic actions, enabling fine-grained multi-label detection in realistic scenes. Video captioning evolved from MSVD Chen & Dolan (2011) and MSR-VTT Xu et al. (2016), which paired short clips with multiple sentences, to dense captioning in untrimmed videos via ActivityNet Captions Krishna et al. (2017) and procedural datasets like YouCook2 Zhou et al. (2018). Multilingual benchmarks such as VATEX Wang et al. (2019) expanded the task to cross-lingual settings with high-quality parallel annotations. In VQA, static image datasets like VQA v2.0 Goyal et al. (2017) and CLEVR Johnson et al. (2017) laid the foundation for compositional reasoning, while TGIF-QA Jang et al. (2017), MovieQA Tapaswi et al. (2016), and TVQA Lei et al. (2018) introduced temporal and multimodal reasoning in video. Recent efforts such as NExT-QA Xiao et al. (2021) and CLEVRER Yi et al. (2020) focus on causal inference and counterfactual reasoning. However, existing datasets are constrained to terrestrial environments and lack domain-specific complexities inherent to space-based activities. To bridge this gap, we introduce the first benchmark for video captioning and VQA in microgravity, enabling the evaluation of vision-language models in long-horizon, space-relevant scenarios.
@@
Fine-Grained Human Action Recognition. Fine-grained video-based action recognition Gritsenko et al. (2023); Chung et al. (2021) aims to detect fine-grained, indivisible human actions in both single- and multi-person videos. Unlike standard clip-level recognition, atomic action localization requires frame-level, multi-label classification with spatio-temporal bounding box predictions. To tackle this, CNN-based Wang et al. (2023a); Feichtenhofer (2020); Feichtenhofer et al. (2019) and transformer-based Ryali et al. (2023); Wang et al. (2023a); Li et al. (2022); Wang et al. (2023b; 2022); Peng et al. (2022); Gritsenko et al. (2023) models have been adapted by adding multi-label heads, bounding box regressors, and region-of-interest modules. Notably, Ryali et al.Ryali et al. (2023) introduced a hierarchical vision transformer balancing accuracy and efficiency, while Wanget al. Wang et al. (2023a) proposed a dual-masked autoencoder for improved video pretraining. In this work, we evaluate such HAR methods for the first time in microgravity scenarios.
@@ 3 Collection Methodology
Filtering. Video clips undergo automatic filtering based on person detection and scene-transition analysis. Specifically, we employed YOLOv11 Khanam & Hussain (2024) for human detection and PySceneDetect Castellano to identify abrupt scene changes, discarding clips with insufficient human-action content or disrupted temporal continuity.
Bounding-box Annotation. Person bounding boxes are automatically annotated using YOLOv11 Khanam & Hussain (2024) detection combined with BoT-SORT tracking Aharon et al. (2022). To enhance annotation accuracy, adaptive strategies were employed by assessing video motion intensity using sparse optical flow methods Jeannin & Divakaran (2001); Ali (2013); Szeliski (2010). Annotation parameters were dynamically adjusted accordingly to optimize computational efficiency and annotation precision. The resulting structured annotations include bounding-box coordinates, unique identities, and detection confidence scores.
@@
[…]
@@
For Video Question Answering (VQA), we adopted a structured approach based on Heilman & Smith (2009), formulating diverse questions using standard interrogative forms. Based on the complexity of the video captions, Large Language Models (LLMs) generate a proportional number of candidate question-answer pairs, explicitly filtering out cases with missing or uncertain information. These candidates are then ranked by the LLMs according to logical consistency, linguistic fluency, semantic relevance, and informational value. Subsequently, multiple rounds of human-assisted review and prompt adjustment are employed to eliminate any hallucinated or fabricated content produced by the LLMs. Finally, a concise set of six diverse and content-rich QA pairs is selected for each video, ensuring comprehensive coverage from broad contextual understanding to detailed actions.
@@ 4 Dataset Composition
Fine-Grained Human Action Recognition Subset: Contains fine-grained, multi-label annotations for spatiotemporal human action detection, comprising $4,759$ manually annotated three-second video clips from authentic microgravity environments. Each clip includes annotations for up to five distinct actions per detected individual. The $50$ action labels are organized into three categories: Object Manipulation ($4,976$ annotations, $37.60\%$), Person Interaction ($4,288$ annotations, $32.34\%$), and Person Movement ($3,987$ annotations, $30.07\%$). Figure 2(d) illustrates the distribution of these categories in both real and simulated video clips, while Figure 2(e) shows the distribution of person counts per clip, highlighting the diversity of social configurations. Overall, the subset contains $13,261$ action annotations, including $9,610$ from real footage, $3,651$ from simulated sources, and $390,000$ bounding box annotations.
@@
Figure 2 presents representative video clips from the dataset, demonstrating typical annotation examples and highlighting the visual and semantic diversity of the dataset. More details regarding the dataset can be found in the appendix.
@@ 5.1 Benchmark Protocol
Data Split. We partition the dataset into training, validation, and test subsets in a 7:1:2 ratio. Row‐level distribution is as follows: Training set: 9,266 records (69.93% of 13,251); Validation set: 1,329 records (10.03% of 13,251); Test set: 2,656 records (20.04% of 13,251). Video‐level distribution is as follows: Training set: 3,331 videos (69.99% of 4,759); Validation set: 475 videos (9.98% of 4,759); Test set: 953 videos (20.03% of 4,759).
Baselines for Fine-Grained HAR in Microgravity Scenarios. For fine-grained action recognition, we evaluate well-established baselines from video-based human action recognition, including transformer-based models (MViTv1 Fan et al. (2021), MViTv2 Li et al. (2022)) and CNN-based models (I3D Carreira & Zisserman (2017), SlowFast Feichtenhofer et al. (2019), X3D Feichtenhofer (2020), C2D Feichtenhofer et al. (2019)). These widely used architectures cover both paradigms and allow us to assess generalization to the unique spatiotemporal dynamics of microgravity. All models are initialized with Kinetics400 Kay et al. (2017) pretrained weights for better convergence.
Baseline methods for Video Captioning and QA in Microgravity Scenarios. For video captioning and question answering in microgravity, we evaluate strong baselines, including open-source models (VideoChatGPT Li et al. (2023), mPLUG-Owl-3 Ye et al. (2025), LLaVA-Next Li et al. (2024), VideoLLaVa Lin et al. (2024), Qwen-2.5-VL Bai et al. (2025), InternVideo Wang et al. (2022)) and closed-source models (GPT-4o Hurst et al. (2024), Gemini 1.5 Pro Reid et al. (2024)). Open-source models offer reproducibility via public weights and code, while closed-source models serve as upper-bound references. This mix enables both transparent analysis and comprehensive evaluation of video-language understanding in microgravity.
Cross-domain transfer protocol. To quantify the microgravity-induced domain gap, we use a fixed transfer setup: models pretrained on Kinetics are fine-tuned on AVA Gu et al. (2018) with matched settings, then evaluated zero-shot on MicroG-4M. For terrestrial contrast, we test on JHMDB (Split 1) Jhuang et al. (2013) with the overlapping action set and standard protocol. This isolates domain effects (microgravity vs. Earth) from implementation choices; evaluation details in Sec. 5.
Evaluation Metrics for Fine-Grained HAR. Our evaluation metrics include mAP@0.5, F1 score, recall, and AUROC, all calculated using the macro method. Among these, mAP@0.5 is the primary metric for measuring average detection accuracy per category and thus comprehensively evaluating the model’s action recognition performance in a microgravity environment. Per-class threshold sweeps are provided in the appendix.
[…]